OpenStudy (anonymous):

PLEASE HELP!!!!! MEDAL AND FAN!!!!! 26. Give the center coordinates (x,y) of the circle: (x + 3)2 + (y - 4)2 = 36 27. Give the radius of the circle: (x + 3)2 + (y - 4)2 = 36 28. Give the center coordinates (x,y) of the circle: (x + 1)2 + (y - 6)2 = 2.25 29. Give the radius of the circle: (x + 1)2 + (y - 6)2 = 2.25

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

this is 26? \[(x+3)^2+(y-4)^2=36\]

OpenStudy (anonymous):

But how do I solve it?

OpenStudy (anonymous):

center of the circle has got to do with the 3 and the 4

OpenStudy (anonymous):

when the center shift left it becomes (x+h)^2 and (x-h)^2 when it shifts right, same with the y axis

OpenStudy (anonymous):

wait would the answer to number 26 be -3,4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now for the radius you know the circle comes in this package \[(x+3)^2+(y-4)^2=36\]which is equivalent to \[(x+h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

Would it be 6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now use the same method for 28 & 29

OpenStudy (anonymous):

So 28 is -6,1

OpenStudy (anonymous):

And 29 is 1.5

OpenStudy (anonymous):

Is tha correct?

OpenStudy (anonymous):

*that

OpenStudy (anonymous):

no and yes

OpenStudy (anonymous):

you swapped the x and y position for 28

OpenStudy (anonymous):

So it's 1,-6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-1,6

OpenStudy (anonymous):

Oh okay thanks for you help.

OpenStudy (anonymous):

not -6,1

OpenStudy (anonymous):

yw